Sol Summary – November 2nd


Summary Title: Boomer 1 & 2 go on EVA
Author’s name: Guy Murphy
Mission Status: Active.
Sol Activity Summary: The morning was spent attending to hab maintenance, cleaning tasks. During the afternoon, two crew members (Boomer 1 and 2) undertook an EVA to continue collecting micrometeorite samples. 3 meals were prepared today, including toasted sandwiches with Spam & Polish pickels.
Look Ahead Plan: Sunday 3 November (Sol 7) is a rostered day of rest for the crew, with no EVA’s scheduled. We will resume our normal work schedule on Monday 4 November.
Anomalies in work: None.
Weather: Clear skies, no precipitation. Below freezing overnight.
Crew Physical Status: All in good health.
EVA:Afternoon EVA completed successfully, no problems encountered.
Reports to be filed: Sol Summary, EVA Report, Journalists Report, Science Report, Operations Reports, Photos.
Support Requested: No further support requested. The crew is enjoying the new dietary items that were provided to us in yesterday’s delivery.

Copyright © The Mars Society. All rights reserved. | Main Site